# From Champagne to Confined Polymer: Natural and Artificial Bubble Nucleation

### v2

@@ IV Conclusion
Our experimental results also show that bubble growth is directly related to the imposed temperature before decompression, i.e., the higher the temperature, the greater the likelihood of nucleating and growing a bubble. Counter-intuitively, however, the number of nuclei is inversely proportional to the imposed temperature. This may be related to the interaction between the PVB polymer and the confining glass lites. When heated, the polymer chains relax and can even reptate relative to one another, thus allowing the polymer to flow and better conform to the confining glass lites. Probably, the better flow capabilities at elevated temperatures improves air dissolution, thereby diminishing the number of nucleation sites available to grow a bubble.
The nucleation of bubbles seems to be always of the heterogenous kind, i.e., bubbles grow at a polymer/glass interface. This is not to say that homogenous(bulk) nucleation is not possible. However, our accumulated body of experimental work highlights that the former is more likely than the latter. Our post-decompression bubbles grow with a morphology linked to the initial gas nuclei. When the latter is sufficiently small and well-separated from other neighboring bubbles, the bubble takes on a tear-like circular shape. If the nuclei are rather large, the bubble grows with a contour shape that initially mimics the original nuclei shape. Even in the case of circular bubbles, anisotropy can ensue if bubbles become close to one another.
Elegantly, the nucleation mechanism can be subdivided in similar fashion to Champagne’s into natural and artificial nucleation. Natural nucleation consists in bubbles nucleating and growing at a trapped dust speckle or fiber. Artifical nucleation involves bubbles forming at an imperfection in the glass, such as a pit or crenel. We have provided evidence for both. Bubbles growing from natural nucleation exhibit a constant global growth rate, whereas those growing from artificial nucleation experience a growth rate linked to their local conditions, in particular the size of the defect itself.
@@ Appendix A Nitrogen sorption
Transient PVB mass during a sorption experiment for $N_{2}$ (20C) as a function of the applied pressure (absolute). The inset presents the sorption isotherm ($\square$), i.e., the steady uptake of mass $m_{a}$ (with buoyancy correction) with respect to the reference mass $m_{ref}$ (degassed PVB mass).    Henry slope, i.e., the solubility of $N_{2}$.
@@
When comparing the PVB blends, we find the solubility constants to be of comparable magnitude but a large disparity was observed for the average diffusion coefficient. The latter finding is hardly surprising since the glassy state highly constraints the polymer chains, thus hindering the diffusion of molecules. Surprisingly, despite our PVB blend including three monomer units and a plasticizer molecule, the order of magnitude of the solubility constant ($\sim 10^{-5}$ kg kg-1 bar-1), diffusion coefficients ($\sim$10-11m2s-1), and the activation energy for diffusion (20.6-41.9 kJ mol-1, see table 3), were all in the range observed for N2 in the homopolymers of polysioprene or polybutadiene. This highlighted to us that since $N_{2}$ is an inert gas, it rarely interacts with the polymer matrix, thus the similarity of its transport properties across rubber-like materials with a chemistry comprised mainly by carbon, hydrogen, and oxygen. The only quantity extracted from gravimetric analysis that did not fit into this comparison framework was the sign of the heat of solution. Indeed, our gravimetric analysis suggested that the solubility of N2 seemingly decreased when the temperature was raised from 20 to 40°C (See Table 2, runs A,B), which yielded $\Delta H_{s,a}$ $<0$. We believe this result to be unreliable for the following reasons.
Table 1: Comparison of N2 solubility and diffusivity between RB41 and rubbery polymers
@@
First, the experimental resolution of our gravimetric analysis was not sufficient to draw any quantitative conclusion. An additional sorption analysis at 20°C (run C in Table 2) revealed a measurement variability of about 27% (runs A,C), which is larger than the perceived difference with temperature of 12% (runs A,B). Therefore, the measured solubilities at 20 and 40°C are unfortunately in the uncertainty range of the experiment.
Then, it contradicts thermodynamic expectations when considering that nitrogen is incondensable at room conditions. From a theoretical point of view, if we take the solubility process as condensation at the surface followed by mixing in the bulk, then $\Delta H_{s}=\Delta H_{cond}+\Delta H_{1}$, where the former is the heat of condensation (negative in sign) and the latter is the heat of mixing (positive in sign)[19, 32, 20, 21]. For incondensable nitrogen at room conditions, it holds that $\Delta H_{cond}\ll\Delta H_{1}$, and evidently $\Delta H_{s,a}$ $>0$.
Last, our bubble experiments also suggested $\Delta H_{s,a}$ $>0$. Indeed, when using dry PVB, we did not observe any bubble nucleation post decompression, despite extra air being dissolved in the PVB bulk from the anisotropic bubbles that initially populate our glass samples.
@@ 7 $1\times 10^{-10}$
In this paper, we took a middle of the road value of 5 kJ mol-1. A sensitivity analysis over the 1-10 kJ mol-1 range across our experimental conditions reveals that the air oversaturation driver is bounded by $0.51\leq$ $\color[rgb]{1,0,0}{f_{a}}$ $\leq 1.7$, i.e., air is around the saturation point in our experiments regardless of the precise value of $\color[rgb]{1,0,0}{\Delta H_{s,a}}$.

### v3

@@ IV Conclusion
Our experimental results also show that bubble growth is directly related to the imposed temperature before decompression, i.e., the higher the temperature, the greater the likelihood of nucleating and growing a bubble. Counter-intuitively, however, the number of nuclei is negatively correlated with the imposed temperature. This may be related to the interaction between the PVB polymer and the confining glass slides. When heated, the polymer chains relax and can even reptate relative to one another, thus allowing the polymer to flow and better conform to the confining glass slides. The better flow capabilities at elevated temperatures probably improves air dissolution, thereby diminishing the number of nucleation sites available to grow a bubble.
The nucleation of bubbles seems to always be of the heterogenous kind, i.e., bubbles grow at a polymer/glass interface. This is not to say that homogenous (bulk) nucleation is not possible. However, our accumulated body of experimental work highlights that the former is more likely than the latter. Our post-decompression bubbles grow with a morphology linked to the initial gas nuclei. When the latter is sufficiently small and well-separated from other neighboring bubbles, the bubble takes on a circular shape. If the nuclei are rather large, the bubble grows with a contour shape that initially mimics the original nuclei shape because the contact line of the bubble is always pinned. Even in the case of circular bubbles, anisotropy can ensue if bubbles become close to one another.
Elegantly, the nucleation mechanism can be subdivided in a similar fashion to champagne’s into natural and artificial nucleation. Natural nucleation consists of bubbles nucleating and growing at a trapped dust speckle or fiber. Artifical nucleation involves bubbles forming at an imperfection in the glass, such as a pit or crenel. We have provided evidence for both. Bubbles growing from natural nucleation exhibit a constant global growth rate, whereas those growing from artificial nucleation experience a growth rate linked to their local conditions, in particular, the size of the defect itself.
@@ Appendix A Nitrogen sorption
Transient PVB mass during a sorption experiment for $N_{2}$ (20C) as a function of the applied pressure (absolute). The inset presents the sorption isotherm ($\square$), i.e., the steady uptake of mass $m_{a}$ (with buoyancy correction) with respect to the reference mass $m_{ref}$ (degassed PVB mass).    Henry slope, i.e., the solubility constant of $N_{2}$.
@@
When comparing the PVB blends, we find the solubility constants to be of comparable magnitude, but a large disparity was observed for the average diffusion coefficient. The latter finding is hardly surprising since the glassy state highly constrains the polymer chains, thus hindering the diffusion of molecules. Surprisingly, despite our PVB blend including three monomer units and a plasticizer molecule, the order of magnitude of the solubility constant ($\sim 10^{-5}$ kg kg-1 bar-1), diffusion coefficients ($\sim$10-11m2s-1), and the activation energy for diffusion (20.6-41.9 kJ mol-1, see table 3), were all in the range observed for N2 in the homopolymers of polysioprene or polybutadiene. This highlighted to us that since $N_{2}$ is an inert gas, it rarely interacts with the polymer matrix, thus the similarity of its transport properties across rubber-like materials with a chemistry comprised mainly by carbon, hydrogen, and oxygen. The only quantity extracted from gravimetric analysis that did not fit into this comparison framework was the sign of the heat of solution. Indeed, our gravimetric analysis suggested that the solubility constant of N2 seemingly decreased when the temperature was raised from 20 to 40°C (See Table 2, runs A and B), which yielded $\Delta H_{s,a}$ <0<0< 0. We believe this result to be unreliable for the following reasons.
Table 1: Comparison of N2 solubility constant and diffusivity between RB41 and rubbery polymers
@@
First, the experimental resolution of our gravimetric analysis was not sufficient to draw any quantitative conclusion. An additional sorption analysis at 20°C (run C in Table 2) revealed a measurement variability of about 27% (runs A and C), which is larger than the perceived difference with temperature of 12% (runs A and B). Therefore, the measured solubility constants at 20 and 40°C are unfortunately in the uncertainty range of the experiment.
Then, it contradicts thermodynamic expectations when considering that nitrogen is incondensable at room conditions. From a theoretical point of view, if we take the solubility process as condensation at the surface followed by mixing in the bulk, then $\Delta H_{s}=\Delta H_{cond}+\Delta H_{1}$, where the former is the heat of condensation (negative in sign) and the latter is the heat of mixing (positive in sign)[19, 32, 20, 21]. For incondensable nitrogen at room conditions, it holds that $\Delta H_{cond}\ll\Delta H_{1}$, and evidently $\Delta H_{s,a}$ >0>0> 0.
Last, our bubble experiments also suggested $\Delta H_{s,a}$ >0>0> 0. Indeed, when using dry PVB, we did not observe any bubble nucleation post decompression, despite extra air being dissolved in the PVB bulk from the anisotropic bubbles that initially populate our glass samples.
@@ 7 $1\times 10^{-10}$
In this paper, we took a middle of the road value of 5 kJ mol-1. A sensitivity analysis over the 1-10 kJ mol-1 range across our experimental conditions reveals that the air relative saturation is bounded by 0.510.51$\leq$0.51 ≤ $\color[rgb]{1,0,0}{f_{a}}$ 1.7$\leq$ 1.7≤ 1.7, i.e., air is around the saturation point in our experiments regardless of the precise value of $\color[rgb]{1,0,0}{\Delta H_{s,a}}$.